MaxiCare Corporation, a not-for-profit organization, specializes in health care for senior citizens. Management is considering whether to expand operations by opening a new chain of care centers in the inner city of large metropolitan areas. For a new facility, initial cash outlays for lease, renovations, net working capital, training, and other costs are expected to be about $16 million. The corporation expects the cash inflows of each new facility in its first year of operation to equal the initial investment outlay for the facility. Net cash inflows are expected to increase to $1.5 million in each of years 2 and 3; $2.5 million in year 4; and $3.0 million in each of years 5 through 10. The lease agreement for the facility will expire at the end of year 10, and MaxiCare expects the cost to close a facility will pretty much exhaust all cash proceeds from the disposal. Cost of capital for Maxicare is estimated as 12%. Assume that all cash flows occur at year end Required 1. Compute (using the built-in NPV function in Excel) the net present value (NPV) the proposed investment. (Negative amount should be indicated by a minus sign. Enter your answer in whole dollars, not in millions, rounded to nearest whole dollar.) 2. Compute (using the built-in IRR function in Excel) the internal rate of return (IRR) for the proposed investment. (Round your final answer 2 decimal places. (i.e. .1234-12.34%)) 3. What is the breakeven selling price for this investment, that is, the price that would yield an NPV of $0? (Use the Goal Seek function in Excel to determine the breakeven selling price. The following online tutorial may be helpful to you: Goal Seek Tutorial.) (Enter your answer in whole dollars, not in millions, rounded to nearest whole dollar.)This is a numeric cell, so please enter numbers only. 1. Estimated NPV 2. Estimated IRR 3. Breakeven selling price 0 0
Answer #1
1 & 2
Year Cash Flow PV @12% Present value
0 -16000000 1 $ -1,60,00,000.00
1 0 0.89285714 $                       -  
2 1500000 0.79719388 $     11,95,790.82
3 1500000 0.71178025 $     10,67,670.37
4 2500000 0.63551808 $     15,88,795.20
5 3000000 0.56742686 $     17,02,280.57
6 3000000 0.50663112 $     15,19,893.36
7 3000000 0.45234922 $     13,57,047.65
8 3000000 0.40388323 $     12,11,649.68
9 3000000 0.36061002 $     10,81,830.07
10 3000000 0.32197324 $       9,65,919.71
NPV $         -43,09,123
IRR 6.28%
3)
Break -even selling price = -11,690,877
Year Cash Flow
0 $         -1,16,90,877.43
1 0
2 1500000
3 1500000
4 2500000
5 3000000
6 3000000
7 3000000
8 3000000
9 3000000
10 3000000
$                               -  
1 & 2
Year Cash Flow PV @12% Present value
0 -16000000 =1/1.12^A13 =B13*C13
1 0 =1/1.12^A14 =B14*C14
2 1500000 =1/1.12^A15 =B15*C15
3 1500000 =1/1.12^A16 =B16*C16
4 2500000 =1/1.12^A17 =B17*C17
5 3000000 =1/1.12^A18 =B18*C18
6 3000000 =1/1.12^A19 =B19*C19
7 3000000 =1/1.12^A20 =B20*C20
8 3000000 =1/1.12^A21 =B21*C21
9 3000000 =1/1.12^A22 =B22*C22
10 3000000 =1/1.12^A23 =B23*C23
NPV =SUM(D13:D23)
IRR =IRR(B13:B23)
3)
Break -even selling price = -11,690,877
Year Cash Flow
0 -11690877.4294121
1 0
2 1500000
3 1500000
4 2500000
5 3000000
6 3000000
7 3000000
8 3000000
9 3000000
10 3000000
=NPV(12%,B31:B40)+B30